Foreign Exchange Risk and the Cost of Borrowing Swiss Francs.  The chapter demonstrated that a firm borrowing in a foreign currency could potentially end up paying a very different effective rate of interest than what it expected. Using the same baseline values of a debt principal of SF1.40 ​million, a​ one-year period, an initial spot rate of SF1.5500​/$, a 4.661​% cost of​ debt, and a 40​% tax​ rate, what is the effective​ after-tax cost of debt for one year for a U.S.​ dollar-based company if the exchange rate at the end of the period​ was SF1.4470​/$, what is the effective​ after-tax cost of​ debt?
Answer #1

Spot rate $1=   1.5500
Debt Principal = SF   1400000
Equivalent dollar Amount = 1400000/1.55=   $903,225.81
cost of Debt =   4.661%
Interest on debt = 1400000*4.661%=   65,254.00
Exchange rate 1$=   $1.4470
interest Equivalent to Dollars = 65254/1.4470=   45,096.06
tax rate =   40%
  
Effective after tax cost of debt = (dollar amount of interest*(1-tax rate))/Dollar amount of principal  
(45096.06*(1-40%))/903225.81  
0.0300   or 3%
  
So effective after tax cost of debt is 3%